# ID-centric Pre-training for Recommendation

### v1

@@ Abstract.
Pre-trained Recommendation, Multi-domain Recommendation
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 2.2. Overview
The framework of IDP is illustrated in Fig.3. Our IDP mainly includes three parts: (1) ID-based sequential model pre-training (2) Cross-domain ID Matcher (3) downstream fine-tuning. Specifically, we first pre-train an ID-based sequential model on a multi-domain pre-training dataset. Subsequently, in order to model the behavior-aware item modality correlations in recommendation, we train a Cross-domain ID Matcher (CDIM) with collaborative signals. With the help of CDIM, we can leverage textual information as a bridge to connect downstream items with pre-trained item embeddings. After that, we utilize the selected pre-trained item embeddings obtained by CDIM to generate ID embeddings for the downstream domain. Lastly, we fine-tune both the pre-trained model and ID-generated ID embeddings for the new domain task. Through this framework, our IDP could handle multiple new domains while taking advantage of powerful and efficient ID-based sequential modeling.
Figure 3. Overall architecture of our IDP.
@@ 2.3. Pre-trained ID-based Sequential Model
In this section, we introduce the pre-training stage of our IDP. Sequential recommendation has been widely studied in previous works (Kang and McAuley, 2018; Sun et al., 2019). Following (Zhou et al., 2020; Xie et al., 2022; Hou et al., 2022), we adopt the classical SASRec (Kang and McAuley, 2018) as our base sequential model. Specifically, for a user sequence $s_{u}=\{v_{1}^{u},v_{2}^{u},...,v_{|s_{u}|^{u}}\}$, we denoted the $l$-th self-attentive layer’s input behavior matrix as $\bm{H}_{u}^{l}=\{\bm{h}_{u,1}^{l},\bm{h}_{u,2}^{l},\cdots,\bm{h}^{l}_{u,|s_{u}|}\}$, where $\bm{h}_{u,i}^{l}$ is the $i$-th behavior’s representation of $u$ at the $l$-th layer. Then we get the $l+1$ layer’s behavior matrix $\bm{H}_{u}^{l+1}$ and user representation $\bm{u}_{o}$ (the last behavior’s representation at $L$-th layer) as follows:
(2) $\bm{H}_{u}^{l+1}=\mathrm{Transformer}^{l}(\bm{H}_{u}^{l}),$
$\bm{u}_{o}=f_{seq}(s_{u}|\Theta)=\bm{h}_{u,|s_{u}|}^{L}.$
@@
(3) $\begin{split}L_{o}=-\sum_{(u,v_{i})\in S^{+}}\sum_{(u,v_{j})\in S^{-}}\log\sigma(\bm{u}_{o}^{\top}\bm{v}_{i}-\bm{u}_{o}^{\top}\bm{v}_{j}),\quad u\in U^{S},v\in V^{S}.\end{split}$
@@ 2.4.1. Textual similarity
Actually, measuring the similarity of two items for textual modality is similar to matching two sentences. Previous works directly adopt BERT (Hou et al., 2022) as the textual information encoder, while BERT is not designed for the sentence-matching task. Considering this, we utilize SimCSE (Gao et al., 2021) to encode the textual information. SimCSE employs contrastive learning to enable the BERT model (Devlin et al., 2019) to generate high-quality sentence embeddings, resulting in excellent performance on the sentence-matching task. Specifically, given an item $i$ and corresponding textual information $c_{v_{i}}=\{w_{1},w_{2},...,{w_{|c_{v_{i}}|}}\}$, we first concatenate a special token $[CLS]$ with $c_{v_{i}}$, then we feed the concatenated text into SimCSE as:
(4) $\begin{split}t_{i}=SimCSE(\{[CLS],w_{1},w_{2},...,{w_{|c_{v_{i}}|}}\}).\end{split}$
$t_{i}$ is the representation of $[CLS]$ token from the last layer of $SimCSE$.
Furthermore, considering that SimCSE is training on the open-accessed corpus, and may exist a gap between the recommendation corpus. We further fine-tune SimCSE on recommendation data with a contrastive learning task. Similar to SimCSE, we simply feed the same input textual information to the encoder twice. In this way, we get two vectors $t_{i}$ and $t_{i}^{{}^{\prime}}$, which are encoded by two randomly dropped encoders (the encoder has dropout operation), respectively. We treat $t_{i}$ and $t_{i}^{{}^{\prime}}$ as a positive pair. The contrastive learning task can be formulated as:
(5) $\begin{split}l_{v_{i}}^{text}=\frac{e^{sim(t_{i},t_{i}^{{}^{\prime}})/\tau}}{e^{sim(t_{i},t_{i}^{{}^{\prime}})/\tau}+\sum_{k}^{N-1}e^{sim(t_{i},t_{k})/\tau}}.\end{split}$
@@
2.4.2. Behavior-involved Fine-tuning
As discussed before, behavioral information is crucial and should also be taken into account when measuring similarity in recommendation. Inspired by the contrastive learning task of $SimCSE$, we utilize collaborative signals to help model learning the behavior similarity in the recommendation scenario. Specifically, we first train a recommendation model on pre-training datasets. In this work, we directly use the pre-trained sequential model in the last step. Subsequently, for each item $v_{i}$ in pre-training datasets, we obtain its most similar items $V_{s}^{v_{i}}=(v_{s,1}^{v_{i}},v_{s,2}^{v_{i}}...,v_{s,k}^{v_{i}})$ in pre-training datasets by embedding similarity. Last we conduct contrastive learning to fine-tune SimCSE. Here, We regard $<v_{i},v_{j}>,v_{j}\in V_{s}^{v_{i}}$ as a positive pair and $<v_{i},v_{k}>,v_{k}\notin V_{s}^{v_{i}}$ as a negative pair. For an item $v_{i}$, the contrastive learning task can be formulated as:
(6) $\begin{split}l_{v_{i}}^{behavior}=\sum_{v_{j}\in V_{s}^{v_{i}}}\frac{e^{sim(t_{i},t_{j})/\tau}}{e^{sim(t_{i},t_{i}^{{}^{\prime}})/\tau}+\sum_{k}^{N-1}e^{sim(t_{i},t_{k})/\tau}}.\end{split}$
@@ 2.5. Generate ID Embedding in New Domains
With the help of CDIM, we can measure the similarity of items by textual information. Intuitively, items that are similar may display analogous interaction patterns in the recommendation, which is stored in ID embedding. Hence, for an item in the downstream domain, we leverage item’s textual information as a bridge to connect this item with related pre-trained item ID embeddings to generate its ID embedding. This process can be formulated as follows:
(8) $\begin{split}V_{v_{i}^{T}}=TOP([sim(t_{i}^{T},t_{1}^{S}),sim(t_{i}^{T},t_{2}^{S}),...,sim(t_{i}^{T},t_{|V^{S}|}^{S})],m),\end{split}$
@@
(9) $\begin{split}score_{i,j}=sim(t_{i}^{T},t_{j}^{S})/\sum_{k}^{m}{sim(t_{i}^{T},t_{k}^{S})},\\ e_{i}^{T}=\sum_{k}^{m}{score_{i,k}*e_{k}^{S}},\end{split}$
@@ 2.6. Fine-tuning Pre-trained Model
After generating ID embeddings for the target domain. We optimize pre-trained model with the supervision of the target domain’s behavior data. Different from previous methods (Hou et al., 2022) that are bonded with the pre-trained encoder, our IDP is flexible. As we transfer knowledge by ID embedding. Our IDP can be fine-tuned in the following ways: (1) we can directly deploy the model in downstream domains with the generated new ID embedding and pre-trained sequential encoder. (2) we can fine-tune both pre-trained sequential encoder and generated ID embedding on downstream data. (3) We can re-train another encoder (it can be a different architecture from the pre-trained model) and fine-tune generated embedding on downstream data. This process can be formulated as follows:

### v2

@@ Abstract.
Pre-trained Recommendation, ID-based Recommendation, Multi-domain Recommendation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Recommender systems
@@ 1. Introduction
2. Related Works
2.1. Sequential Recommendation
Sequential recommendation models mainly leverage users’ chronological behavior sequences to learn users’ preferences. Early works model sequence patterns by Markov chains (MCs). Shani et al. (Shani et al., 2005) formalized sequential recommendation as a sequential decision problem and adopted Markov Decision Processes (MDPs) to solve it. Later (Rendle et al., 2010) combine the power of MCs and MFs and achieve a promising result. High-order sequential relations also be considered in the following work(He and McAuley, 2016; He et al., 2017a). Recently, with the great progress of deep learning, ID-based sequential models have achieved great success, which project item ID to unique dense embedding. GRu4Rec introduces the GRU network to session-based recommendation for ranking tasks. On the top of GRU4Rec, memory networks(Chen et al., 2018; Huang et al., 2018), hierarchical structures(Quadrana et al., 2017), attention mechanisms (Liu et al., 2018), reinforcement learning (Xin et al., 2020) .etc are introduced. Expected recurrent neural network, other models also employed for sequential recommendation. Caser (Tang and Wang, 2018) regards the user behavior sequence as a one-dimensional image and adopts the Convolutional Neural Network (CNN) to capture sequence patterns.
Inspired by the success of Transformer and BERT (Vaswani et al., 2017; Devlin et al., 2019), attention and transformer-based models are introduced to the recommendation. DIN (Zhou et al., 2018) applies attention mechanisms between the target item and user behavior sequence to capture the user interest. SASRec (Kang and McAuley, 2018) utilizes the single-directional Transformer block to model the sequential. Following BERT(Devlin et al., 2019), BERT4Rec (Sun et al., 2019) adopt a bidirectional Transformer to incorporate user behavior information from both directions. FDSA(Zhang et al., 2019) considers the features of items and designs a feature-level self-attention block to leverage the attribute information about items in the user behavior sequence.
2.2. Pre-training in Recommendation.
Recently, pre-training models have achieved great success in NLP (Devlin et al., 2019; Liu et al., 2019) and CV (He et al., 2021; Radford et al., 2021). They aim to transfer prior knowledge from general large-scale datasets to downstream domains. While in recommendation, since IDs are unaligned, the well-learned unique ID embeddings are challenging to transfer directly to downstream new domains. To utilize the power of pre-trained model, some single-domain pre-trained models (Yao et al., 2021; Wu et al., 2021) have been proposed to improve recommendation accuracy of the inner domain. However, these models lack transferability and are unable to apply to cross-domain recommendation. Some works transfer well-learned knowledge by overlapped users/items (Hu et al., 2018; Tsai, 2018; Zhu et al., [n. d.], 2022; Zang et al., 2022; Yuan et al., 2020b). While overlapped users/items generally are limited in the real world, thus they are lack of universality. To build a universal model, some works attempt to design a text-based/modality-based pre-trained model (Hou et al., 2022; Ding et al., 2021; Cui et al., 2022). Those works introduce pre-trained language models (PLM) into recommendations and attempt to replace ID-embedding with modality information. Specifically, UniSRec (Hou et al., 2022) utilizes PLM to learn universal item representations across different domains via items’ textual information. However, recent studies show that ID-embedding is still irreplaceable (Gao et al., 2023; Liu et al., 2023; Hou et al., 2023). Different from previous works, we build an ID-centric pre-trained model, which directly transfers pre-trained ID-embedding to downstream domains.
@@ 3. Methodology
Table 1. Mathematical Notations
Symbols Description
$S$, $T$ The pre-training multi-domain dataset and downstream domain
$U$, $V$ The User set of and the Item set
$u$, $v$ The user and item ID
$s_{u}$ The behavior sequence of user $u$
$E,\mathbf{e_{v}}$ The Item ID embedding Matrix and the ID embedding of item $v$
$P,\mathbf{p_{i}}$ The Position Embedding Matrix and the position embedding of index $i$
in behavior sequence.
$\mathbf{e_{u}}$ The encoded user embedding
$\mathbf{h}_{u}^{0}$ The input representation of sequential encoder
$Trm$ The transformer block
$SimCSE$ The encoder of SimCSE
$c_{v}$ the textual information of item $v$
$t_{v}$ the encoded textual representation of encoded item
$m,TOP(\cdot,\cdot)$ The number of selected similar items of Pre-train
domain and the function that finds top-m items
$V_{v^{T}}$ The set of selected top-m similar items of the Pretraning domain
$f_{seq}$ The sequential encoder of pre-training
$f_{seq}^{new}$ The sequential encoder that re-trained on target downstream domain
$\Theta,\Theta^{*}$ The parameters of the pre-trained sequential encoder and the parameters
of fine-tuned sequential encoder
$\vartheta^{*}$ The parameters of re-trained new sequential encoder by target downstream
domain
@@ 3.2. Overview
The framework of IDP is illustrated in Fig.3. Our IDP mainly includes three parts: (1) ID-based sequential model pre-training (2) Cross-domain ID Matcher (3) downstream fine-tuning. Specifically, we first pre-train an ID-based sequential model on a multi-domain pre-training dataset. Subsequently, in order to model the behavior-aware item modality correlations in recommendation, we train a Cross-domain ID Matcher (CDIM) with collaborative signals. With the help of CDIM, we can leverage textual information as a bridge to connect downstream items with pre-trained item embeddings. After that, we utilize the selected pre-trained item embeddings obtained by CDIM to generate ID embeddings for the downstream domain. Lastly, we fine-tune both the pre-trained model and ID-generated ID embeddings for the new domain task. Through this framework, our IDP could handle multiple new domains while taking advantage of powerful and efficient ID-based sequential modeling. We summary the
@@ 3.3. Pre-trained ID-based Sequential Model
In this section, we introduce the pre-training stage of our IDP. Sequential recommendation has been widely studied in previous works (Kang and McAuley, 2018; Sun et al., 2019). Following (Zhou et al., 2020; Xie et al., 2022; Hou et al., 2022), we adopt the classical SASRec (Kang and McAuley, 2018) as our base sequential model. Note that our IDP is flexible and does not depend on specific sequential models, it can be easily applied to other sequential models. We also evaluate it in the experiment section. Specifically, for a user sequence $s_{u}=\{v_{1}^{u},v_{2}^{u},...,v_{|s_{u}|^{u}}\}$, We input it to the sequential model and it can be formulated as:
(2) $\mathbf{e}_{u}=f_{seq}(s_{u}|\Theta,\mathbf{E},\mathbf{P}),where\,f_{seq}\,is\,SASRec,$
@@
(3) $\begin{split}L_{o}=-\sum_{(u,v_{i})\in S^{+}}\sum_{(u,v_{j})\in S^{-}}\log\sigma(\mathbf{e}_{u}^{\top}\mathbf{e}_{v_{i}}-\mathbf{e}_{u}^{\top}\mathbf{e}_{v_{j}}),\quad u\in U^{S},v_{i},v_{j}\in V^{S}.\end{split}$
@@ 3.4.1. Textual similarity
Actually, measuring the similarity of two items for textual modality is similar to matching two sentences. Previous works directly adopt BERT (Hou et al., 2022) as the textual information encoder, while BERT is not designed for the sentence-matching task. Considering this, we utilize SimCSE (Gao et al., 2021) to encode the textual information. SimCSE employs contrastive learning to enable the BERT model (Devlin et al., 2019) to generate high-quality sentence embeddings, resulting in excellent performance on the sentence-matching task. Specifically, given an item $v_{i}$ and corresponding textual information $c_{v_{i}}=\{w_{1},w_{2},...,{w_{|c_{v_{i}}|}}\}$, we first concatenate a special token $[CLS]$ with $c_{v_{i}}$, then we feed the concatenated text into SimCSE as:
(4) $t_{v_{i}}=SimCSE(\{[CLS],w_{1},w_{2},...,{w_{|c_{v_{i}}|}}\}).$
$t_{v_{i}}$ is the representation of $[CLS]$ token from the last layer of $SimCSE$.
Furthermore, considering that SimCSE is training on the open-accessed corpus, and may exist a gap between the recommendation corpus. We further fine-tune SimCSE on recommendation data with a contrastive learning task. Similar to SimCSE, we simply feed the same input textual information to the encoder twice. In this way, we get two vectors $t_{v_{i}}$ and $t_{v_{i}}^{{}^{\prime}}$, which are encoded by two randomly dropped encoders (the encoder has dropout operation), respectively. We treat $t_{v_{i}}$ and $t_{v_{i}}^{{}^{\prime}}$ as a positive pair. The contrastive learning task can be formulated as:
(5) $\begin{split}l_{v_{i}}^{text}=\frac{e^{sim(t_{v_{i}},t_{v_{i}}^{{}^{\prime}})/\tau}}{e^{sim(t_{v_{i}},t_{v_{i}}^{{}^{\prime}})/\tau}+\sum_{k}^{N-1}e^{sim(t_{v_{i}},t_{v_{k}})/\tau}}.\end{split}$
@@
3.4.2. Behavior-involved tuning
As discussed before, behavioral information is crucial and should also be taken into account when measuring similarity in recommendation. Inspired by the contrastive learning task of $SimCSE$, we utilize collaborative signals to help model learning the behavior similarity in the recommendation scenario. Specifically, we first train a recommendation model on multi-domain pre-training datasets. In this work, we directly use the pre-trained sequential model in the last step. Subsequently, for each item $v_{i}$ in pre-training datasets, we obtain its most similar items $V_{s}^{v_{i}}=(v_{s,1}^{v_{i}},v_{s,2}^{v_{i}}...,v_{s,k}^{v_{i}})$ in pre-training datasets by embedding similarity of pre-trained model. Last we conduct contrastive learning to fine-tune SimCSE. Here, We regard $<v_{i},v_{j}>,v_{j}\in V_{s}^{v_{i}}$ as a positive pair and $<v_{i},v_{k}>,v_{k}\notin V_{s}^{v_{i}}$ as a negative pair. For an item $v_{i}$, the contrastive learning task can be formulated as:
(6) $\begin{split}l_{v_{i}}^{behavior}=\sum_{v_{j}\in V_{s}^{v_{i}}}\frac{e^{sim(t_{v_{i}},t_{v_{j}})/\tau}}{e^{sim(t_{v_{i}},t_{v_{j}})/\tau}+\sum_{k}^{N-1}e^{sim(t_{v_{i}},t_{v_{k}})/\tau}}.\end{split}$
@@ 3.5. Generate ID Embedding in New Domains
In this section, we introduce how we use the pre-trained item ID embedding to generate the item ID embedding of downstream scenarios. With the help of CDIM, we can measure the similarity of items by textual information. Intuitively, items that are similar may display analogous interaction patterns in the recommendation, which is stored in ID embedding. Hence, for an item in the downstream domain, we leverage the item’s textual information as a bridge to connect this item with related pre-trained item ID embeddings to generate its ID embedding. This process can be formulated as follows:
(8) $\begin{split}V_{v_{i}^{T}}=TOP([sim(t_{v_{i}}^{T},t_{1}^{S}),sim(t_{i}^{T},t_{2}^{S}),...,sim(t_{i}^{T},t_{|V^{S}|}^{S})],m),\end{split}$
@@
(9) $\begin{split}score_{{v_{i}},{v_{j}}}&=sim(t_{v_{i}}^{T},t_{v_{j}}^{S})/\sum_{k}^{m}{sim(t_{v_{i}}^{T},t_{v_{k}}^{S})},\\ e_{v_{i}}^{T}&=\sum_{k}^{m}{score_{i,k}*e_{v_{k}}^{S}},\end{split}$
@@
3.6. Applied to downstream domain
After generating ID embeddings for the target domain, we can apply the pre-trained model to downstream domains. Different from previous methods that are bonded with a pre-trained encoder, our IDP focuses on transfer the knowledge contained in ID embedding. Thus our IDP is fixable It can be applied in the following ways:
  • We can directly apply the model in downstream domains. In this setting, we directly deploy the generated ID embedding and the pre-trained sequential model to downstream domains without any tuning. This process can be formulated as :
    (10) $\begin{split}\mathbf{e}_{u}^{T}=f_{seq}(s_{u}^{T}|\Theta,\mathbf{E}^{T},\mathbf{P}),\end{split}$
    where the $f_{s}eq$ is the pre-trained sequential model, $\Theta$ is the parameters of the pre-trained model, and it contains the parameters of the embedding layer and sequential encoder .
  • we can fine-tune both pre-trained sequential encoder and generated ID embedding on downstream data. This can be formulated as:
    (11) $\centering\begin{split}\mathbf{e}_{u}^{T}&=f_{seq}(s_{u}^{T}|\Theta^{*},\mathbf{E^{*}}^{T},\mathbf{P}^{*}),\\ \Theta^{*},\mathbf{E^{*}}^{T},\mathbf{P}^{*}&=\arg\max_{\Theta,\mathbf{E}^{T},\mathbf{P}}P(v_{|s_{u}|+1}|s_{u},\Theta),u\in U^{T},v\in V^{T},\end{split}\@add@centering$
    where $\Theta^{*}$ is the fine-tuned parameters.
  • We can re-train another encoder (it can be a different architecture from the pre-trained model) and fine-tune generated embedding on downstream data.
    (12) $\centering\begin{split}\mathbf{e}_{u}^{T}&=f_{seq}^{new}(s_{u}^{T}|\vartheta^{*},\mathbf{E^{*}}^{T},\mathbf{P}^{*}),\\ \vartheta^{*},\mathbf{E^{*}}^{T},\mathbf{P}^{*}&=\arg\max_{\vartheta,\mathbf{E}^{T},\mathbf{P}}P(v_{|s_{u}|+1}|s_{u},\vartheta),u\in U^{T},v\in V^{T},\end{split}\@add@centering$
    where $\vartheta^{*}$ is the parameters of the new sequential model.
@@ 3.7. Fine-tuning Pre-trained Model
After generating ID embeddings for the target domain. We optimize pre-trained model with the supervision of the target domain’s behavior data. Different from previous methods (Hou et al., 2022) that are bonded with the pre-trained encoder, our IDP is flexible as we transfer knowledge by ID embedding. Our IDP can be applied in the following ways:
  • we can directly deploy the model in downstream domains with the generated new ID embedding and pre-trained sequential encoder.
(1) we can directly deploy the model in downstream domains with the generated new ID embedding and pre-trained sequential encoder. (2) we can fine-tune both pre-trained sequential encoder and generated ID embedding on downstream data. (3) We can re-train another encoder (it can be a different architecture from the pre-trained model) and fine-tune generated embedding on downstream data. This process can be formulated as follows:
@@
Algorithm 1 Training process of IDP
1:Pre-training multi-domains user set $\bm{U^{S}}$ and item set $\bm{V^{S}}$;Pre-training dataset $\mathcal{X}^{T}$ ; Pre-trianing Sequential recommendation encoder $\bm{f_{seq}}$; Target downstream domain user set $\bm{U^{T}}$ and target downstream domain user set $\bm{V^{T}}$; Target downstream data $\mathcal{X}^{T}$
2:
3:$\mathbf{\Theta}$, $\mathbf{E}^{S}$, $\mathbf{P}$,$f_{seq}\leftarrow$ Call function: PRE-TRAINING SEQUENTIAL MODEL
4:Call function: TRAINING CDIM
5:$\mathbf{E}^{S}\leftarrow$ Call funciton: GENERATING ID EMBEDDING
6:Fine-tuning by Eq.10,Eq.11 or Eq.12
7:
8:function Pre-training sequential model
9:     Initialize the trainable parameters $\mathbf{\Theta}$ of encoder $f_{seq}(\cdot)$, item embedding matrix $\mathbf{E}$ and position matrix $\mathbf{P}$;
10:     repeat
11:         for batch data in $\mathcal{X}^{S}$  do
12:              $\mathbf{e}_{u}\leftarrow$ encode the user behavior sequence $f_{seq}$(bath data)
13:              $L_{0}\leftarrow$ get the loss with Eq.3
14:              Update $\mathbf{\Theta}$, $\mathbf{E}$ and $\mathbf{P}$ according $L_{0}$
15:         end for
16:     until the trainable parameters have converged.
17:     return The pre-trained $\mathbf{E}$, $\mathbf{P}$, $\mathbf{\Theta}$ and $f_{seq}$
18:end function
19:
20:function Training CDIM
21:     repeat
22:         $L^{text}\leftarrow$ calculate the textual contrastive learning loss with Eq.5
23:         $L^{behavior}\leftarrow$ calculate behavior-involved contrastive learning loss with Eq.6
24:         Ł$\L\leftarrow L^{text}+L^{behavior}$
25:         Update SimCSE by $L$
26:     until the trainable parameters have converged.
27:end function
28:
29:function Generating ID embedding
30:     $V_{v_{i}^{T}}\leftarrow$ get the most similar items in $V^{S}$ by Eq.8
31:     $E^{T}\leftarrow$ generate item embedding with Eq.9
32:end function
3.8. Base Model
In this paper, Following (Zhou et al., 2020; Xie et al., 2022; Hou et al., 2022), we adopt the classical SASRec (Kang and McAuley, 2018) as our base sequential model. SASRec encodes the user behavior sequence by stacking the transformer layers and achieving promising performance. The SASRec consists of three sub-modules: an embedding layer, a multi-head attention module and a position-wise feed-forward Network. Note that our IDP is flexible and does not depend on specific sequential models, it can be easily applied to other sequential models.
3.8.1. Embedding Layer
The embedding layer projects the high-dimensional one-hot item ID to low-dimensional dense embedding by looking up an item embedding table $\mathbf{E}\in\mathbb{R}^{|V|\times d}$. In addition, there is also a position embedding table $\mathbf{P}\in\mathbb{R}^{T\times d}$ that projects the position of items in sequence to dense embedding, where $T$ is the maximum length of the sequence. Given a user behavior sequence $s_{u}=\{v_{1}^{u},v_{2}^{u},...,v_{|s_{u}|}^{u}\}$, we can obtain the input representation of user behavior sequence:
(14) $\begin{split}\mathbf{h}_{u}^{0}&=\{\mathbf{h}_{u,1}^{0},\{\mathbf{h}_{u,2}^{0},...,\mathbf{h}_{u,|s_{u}|}^{0}\},\\ \mathbf{h}_{u,i}^{0}&=\mathbf{e}_{v_{i}}+\mathbf{p}_{i},\end{split}$
where $e_{i}\in E$ is the representation of item $v_{i}$, $p_{i}\in P$ is the representation for position index $i$ and $h_{u,i}^{0}$ is the input representation of $v_{i}$.
3.8.2. Self-attention Block
After the embedding layer, we employ the self-attention block to capture the dependencies between items in the sequence and extract the preference of users. Like transformer(Vaswani et al., 2017). The multi-head self-attention mechanism is adopted to effectively extract selective information in different subspace. Specifically, It first projects the input to $k$ subspace and applies self-attention to each subspace. Then it combines the output of each head by linear projecting the concatenated embedding. this process is formulated as:
$MH(\mathbf{H}^{l})=concat(head_{1};head2;,...;head_{k})W^{O}$
(15) $head_{i}=Attention(\mathbf{H}^{l}W_{i}^{Q},\mathbf{H}^{l}W_{i}^{K},\mathbf{H}^{l}W_{i}^{V})$
where $W_{i}^{Q},W_{i}^{K},W_{i}^{V}\in\mathbb{R}^{d\times d\times\frac{d}{h}}$ and $W^{O}\in\mathbb{R}^{d\times d}$are the learnable project matrix.
In this paper, we adopt the scaled dot-product attention as the attention function:
(16) $Attention(Q,K,V)=softmax(\frac{QK^{\top}}{\sqrt{d/h}})V$
where $K,Q,V$ denote queries, keys and values vector, respectively. And $\sqrt{d/h}$ is the scale factor to avoid large values of inner product.
3.8.3. Point-Wise Feed-Forward Network
The multi-head self-attention mechanism is effective in extracting the relation between items, while the simple linear projection limits the capacity of model. To endow the model with nonlinear interactions between different dimensions, we adopt a Position-wise Feed-Forward Network after the multi-head self-attention mechanism sub-layer, which is formulated as:
$PFFN(\mathbf{H}^{l})=[FFN(\mathbf{h}_{1}^{l})^{\top},...;FFN(\mathbf{h}_{t}^{l})^{\top}]^{\top}$
(17) $FFN(\mathbf{h}_{i}^{l})=GELU(\mathbf{h}_{i}^{l}+b_{1})W_{2}+b2$
where $W_{1},b1,W_{2},b2$ are learnable parameters. $GELU$ is activation function and is widely adopted in other works. (Vaswani et al., 2017; Devlin et al., 2018; Sun et al., 2019)
3.8.4. Stack more Blocks
. Empirically, stacking more layers is benefit for modeling patterns in deep learning. While model generally becomes difficult to train with it goes deeper. Residual connection, dropout and layer normalization module are proved effectively to alleviate this issue. So we apply them on our model:
(18) $LayerNorm(x+Dropout(sbulayer(x)))$
where $sublayer$ is the above self-attention block.
In summary, our sequence-based modeling module is represented as follows:
$\mathbf{H}^{0}=\mathbf{E}+\mathbf{P}$
$\mathbf{H}^{l}=Trm(\mathbf{H}^{l-1})$
$Trm(\mathbf{H}^{l-1})=LayerNorm(\mathbf{A}^{l-1}+Dropout(PPFN(\mathbf{A}^{k-1})))$
(19) $\mathbf{A}^{l-1}=LayerNorm(\mathbf{H}^{l-1},Dropout(MH(\mathbf{H}^{l-1})))$
Input a behavior sequence of a user, our multi-layer self-attention blocks output a series of representations of all step $t$. In here, we adopt the last step to output the sequence as the user’s local preference under this behavior. We denote it as $\mathbf{e_{u}}=\bm{h}_{u,|s_{u}|}^{L}$.